### v1

# LLM2Vec-Gen: Generative Embeddings from Large Language Models

Abstract
LLM-based text embedders typically encode the semantic content of their input. However, embedding tasks require mapping diverse inputs to similar outputs. Typically, this input-output is addressed by training embedding models with paired data using contrastive learning. In this work, we propose a novel self-supervised approach, LLM2Vec-Gen, which adopts a different paradigm: rather than encoding the input, we learn to represent the model’s potential response. Specifically, we add trainable special tokens to the LLM’s vocabulary, append them to input, and optimize them to represent the LLM’s response in a fixed-length sequence. Training is guided by the LLM’s own completion for the query, along with an unsupervised embedding teacher that provides distillation targets. This formulation helps to bridge the input-output gap and transfers LLM capabilities such as safety alignment and reasoning to embedding tasks. Crucially, the LLM backbone remains frozen and training requires only unlabeled queries. LLM2Vec-Gen achieves state-of-the-art self-supervised performance on the Massive Text Embedding Benchmark (MTEB), improving by 9.3% over the best unsupervised embedding teacher. We also observe up to 43.2% reduction in harmful content retrieval and 29.3% improvement in reasoning capabilities for embedding tasks. Finally, the learned embeddings are interpretable and can be decoded into text to reveal their semantic content.
1 Introduction
Text embeddings are central to several natural language processing (NLP) applications such as semantic search, textual similarity, and retrieval-augmented generation (Reimers and Gurevych, 2019; Karpukhin et al., 2020). The success of large language models (LLMs) has led to the development of powerful LLM-based text encoders (Zhang et al., 2025b; Lee et al., 2025; BehnamGhader et al., 2024; Muennighoff et al., 2024), which leverage the web-scale pretraining of LLMs to achieve state-of-the-art results across a wide array of text encoding tasks (MTEB; Muennighoff et al., 2023; Enevoldsen et al., 2025).
Although LLMs have advanced the state-of-the-art for text embedders, they still operate under the same input-centric representation paradigm adopted by BERT-based encoders (Reimers and Gurevych, 2019; Gao et al., 2021), i.e., models are trained to produce representations that capture the semantic content of the input text.111This is typically realized by taking the mean over input token representations produced by the embedding model. This paradigm struggles with the inherent nature of embedding tasks where diverse inputs are mapped to similar outputs in a shared embedding space. For example, in clustering, news articles about the same event written from different perspectives should be grouped together. Consequently, training an encoder typically requires contrastive learning on large-scale curated paired data to bridge this gap (Wang et al., 2024; Lee et al., 2025; Zhang et al., 2025b).
In this work, we adopt a different paradigm for training LLM-based text encoders: rather than encoding the input itself, the model should encode the LLM’s potential response to that input (see Figure 1). This paradigm enables bridging the input-output gap and transfers LLM capabilities such as safety and reasoning into the embedding space. For instance, given a harmful query, traditional input-centric encoders represent the malicious intent of the query itself, whereas approaches adopting this paradigm encode the model’s safe refusal (e.g., “I can’t assist with that”). Similarly, reasoning capabilities which manifest in the LLM’s response rather than the input, can be transferred to the embedding space under this paradigm.
We introduce LLM2Vec-Gen, as an instantiation of this paradigm. LLM2Vec-Gen is a self-supervised framework for training LLM-based text encoders to distill a model’s potential response into a fixed set of latent suffix embeddings. More specifically, given a set of unlabeled queries, we first generate responses using the LLM itself (see Figure 2). Next, we add special trainable tokens to the vocabulary and append them to the query as placeholders for the response. The last layer hidden states of these tokens are then passed through shallow projection layers to shape the response embedding. We train LLM2Vec-Gen with two complementary objectives: (1) response reconstruction, where conditioned on the embeddings, the frozen LLM reconstructs its own response in a separate forward pass, and (2) embedding alignment, where we minimize distance between the generated embedding and an unsupervised teacher’s embedding of the LLM’s response. Crucially, the LLM itself remains frozen and only the special tokens and lightweight projection layers are trained.
We apply LLM2Vec-Gen to differently sized models from Llama-3.x, Qwen-2.5, and Qwen-3 model families, using corresponding unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers. We show that our proposed framework substantially outperforms existing unsupervised and self-supervised approaches on MTEB, closing over 60% of the gap to supervised methods. Notably, training with LLM2Vec-Gen consistently surpasses the off-the-shelf embedding teacher used during training, by as much as 9% for some models, which validates the effectiveness of encoding potential LLM responses. On AdvBench-IR (BehnamGhader et al., 2025), which measures retriever safety when presented with adversarial queries, models trained with LLM2Vec-Gen demonstrate superior safety, improving by 9.8 points over LLM2Vec-Qwen-3-8B. On BRIGHT (Su et al., 2025), a reasoning-intensive retrieval benchmark, LLM2Vec-Gen achieves up to 29.3% improvement over input-centric baselines, demonstrating that reasoning capabilities transfer to the embedding space. Finally, we show that our special token embeddings are interpretable and can be decoded back into text, revealing the semantic content they capture.222Our code, data, and pre-trained models are publicly available at https://github.com/McGill-NLP/llm2vec-gen
2 Related work
Recent work on embedding models has increasingly focused on repurposing decoder-only LLMs for text embedding tasks, leveraging their massive web-scale pre-training to improve performance on embedding benchmarks (Wang et al., 2024; Lee et al., 2025). Approaches such as GritLM (Muennighoff et al., 2024), which unify generation and representation within a single model, achieve strong performance on the MTEB benchmark (Muennighoff et al., 2023), but still rely on supervised contrastive learning (Khosla et al., 2020) and large-scale curated labeled datasets to align representations.
To mitigate the need for labeled data, several works have explored unsupervised embedding approaches (Zhang et al., 2025a; Lin et al., 2025; Thirukovalluru and Dhingra, 2025; Jiang et al., 2024). LLM2Vec (BehnamGhader et al., 2024) shows that simple modifications like bidirectional attention and masked next-token prediction, combined with unsupervised SimCSE (Gao et al., 2021), can transform decoder-only LLMs into strong encoders. Similarly, Echo Embeddings (Springer et al., 2025), provides strong unsupervised performance, leveraging a zero-shot repetition strategy, where a model processes an “echoed” version of the input to better aggregate contextual information.
While input-centric approaches improve embedding quality, they often struggle with the lexical and conceptual gap between queries and documents. HyDE (Gao et al., 2023), one of the earliest approach under the “generate-then-encode” paradigm, demonstrates that encoding a hypothetical LLM-generated answer can yield superior retrieval performance compared to encoding the query itself. However, it incurs significant inference latency as it requires generating (long) documents during inference.
More recently, several approaches focused on internalizing this generative foresight into the embedding space itself. InBedder (Peng et al., 2024) treats the instruction as a question about the input and derives embeddings from the first hidden state of the model’s generation, demonstrating that generation-derived representations can outperform prompt-based ones. GIRCSE (Tsai et al., 2026) generates soft tokens autoregressively and refines them with stepwise contrastive loss, achieving strong performance through iterative semantic refinement. Both methods rely on supervised data: InBedder requires abstractive QA pairs, while GIRCSE uses contrastive data with hard negatives. CLaRa (He et al., 2025) compresses documents into learnable memory tokens for RAG efficiency. It jointly optimizes retrieval and generation via next-token prediction loss. LLM2Vec-Gen goes beyond token-level alignment using embedding distillation to ensure embeddings semantically represent potential responses while preserving the efficiency of single-vector retrieval.
Joint Embedding Predictive Architectures (JEPAs) advocate learning by predicting in representation space rather than reconstructing raw inputs (Sobal et al., 2022). Recent work extends this philosophy to language: LLM-JEPA introduces JEPA-style objectives for LLM pretraining and fine-tuning that replace or augment next-token prediction with embedding-space prediction (Huang et al., 2025). Adopting this perspective, our alignment objective shares similarities with the motivation behind JEPA: instead of matching token-level outputs, LLM2Vec-Gen predicts a target representation corresponding to the model’s likely response (via distillation from a teacher encoder), while our reconstruction objective additionally keeps the learned representations grounded in natural language and decodable.
3 LLM2Vec-Gen
The goal of LLM2Vec-Gen is to produce embeddings which represent the output the LLM would have generated for a given query, without actually generating the response at inference time. Our approach consists of the following steps (illustrated in Figure 2):
Formally, let be a pre-trained LLM and be a large corpus of queries. For every , we generate a response using . We introduce two types of new special tokens to ’s vocabulary: thought tokens () and compression tokens (). The thought tokens act as an intermediate computational buffer, while the compression tokens’ () role is to capture the semantic content of the response. For a given query , we append the thought and compression tokens and pass the combined sequence through the LLM to obtain the last layer hidden representations of the compression tokens (): , which are subsequently used for computing two objectives:
Reconstruction objective.
The first objective ensures that the compression tokens () retain sufficient information to reconstruct the target response , i.e., we treat as a compressed representation that must preserve the content of . Concretely, we process them with a lightweight projection layer, , and feed the resulting representations as a set of soft prompts to the LLM for a second forward pass. Conditioned on this soft prompt, we train the model to reconstruct the target response via standard next-token prediction:
Embedding alignment objective.
Optimizing only the reconstruction objective might result in the soft prompts encoding the query rather than a compressed version of the target response, the goal of LLM2Vec-Gen. Consequently, we introduce a second objective to overcome this issue. For the embedding alignment objective, we utilize the encoder teacher model , which provides a target embedding . The projected representations are processed by another projection layer followed by mean pooling, , to minimize the following alignment loss:
We note that this objective is conceptually related to JEPA (Sobal et al., 2022), in that it predicts a semantic target in embedding space. Unlike JEPA, our target embedding comes from an external teacher encoder, whereas JEPA typically predicts representations produced by a learned internal target encoder.
Training and inference.
The final loss combines both objectives: . Throughout training, only the special tokens and the two MLPs are updated, while the LLM remains frozen throughout. Notably, at inference time, LLM2Vec-Gen requires only a single forward pass: we append the special tokens to the input, extract the hidden states of the compression tokens, and apply and to obtain the embedding .
4 Experiments
We evaluate LLM2Vec-Gen along three axes: general text embedding, malicious retrieval, and reasoning-intensive retrieval. After describing our setup (Sections 4.1, 4.2 and 4.3), we present results showing that LLM2Vec-Gen achieves state-of-the-art self-supervised performance and transfers LLM capabilities such as safety alignment and reasoning into the embedding space (Sections 4.4 and 4.5).
4.1 Experimental setup
Model families.
We apply LLM2Vec-Gen to decoder-only LLMs from three model families: Qwen-3 (0.6B, 1.7B, 4B, 8B; Yang et al., 2025), Qwen-2.5 (0.5B, 1.5B, 3B, 7B; Qwen et al., 2025), Llama-3.2 (1B, 3B; Meta, 2024b), and Llama-3.1 (8B; Meta, 2024a). For all models, we use 10 thought tokens () and 10 compression tokens () unless mentioned otherwise.
Encoder teacher.
This choice is guided by two requirements: 1) The teacher should share the same underlying LLM, ensuring compatible representation spaces. 2) It should have been trained with an unsupervised objective, encouraging faithful representation of the input, as opposed to contrastive objectives, which may collapse semantically distinct texts relevant to the same query (Fu et al., 2022). Hence, we use unsupervised LLM2Vec (BehnamGhader et al., 2024) models as embedding teachers.
Training.
LLM2Vec-Gen requires only an unlabeled corpus of user queries. We use single-turn questions from the Tulu instruction-following dataset (Lambert et al., 2025), sampling 160K queries from its training set. We note that the ground-truth responses from the Tulu dataset are not used for training, instead the model is trained using its own generations. We provide sample responses from different models alongside the Tulu answers in Appendix B. We train for one epoch with a batch size of 32. Training an 8B model takes approximately 3.5 hours on 2 H100 GPUs. Additional training details, including hyperparameters and compute requirements, are provided in Appendix A.
4.2 Evaluation
We evaluate LLM2Vec-Gen performance along three axes: (1) for general text embeddings, (2) for malicious retrieval, and (3) for reasoning-intensive retrieval.
General text embedding.
We evaluate LLM2Vec-Gen on MTEB(eng, v2) (Enevoldsen et al., 2025), a refined version of the Massive Text Embedding Benchmark (Muennighoff et al., 2023). MTEB(eng, v2) contains 41 tasks from seven categories: bitext mining, classification, clustering, pair classification, reranking, retrieval, and semantic textual similarity (STS). Additionally, we construct MTEB-Lite, a subset of 10 tasks that preserves the category distribution of the full benchmark (see Appendix A). We use MTEB-Lite for ablations and analysis. We report the average score across all tasks as the primary metric.
Malicious retrieval.
We additionally evaluate on AdvBench-IR (BehnamGhader et al., 2025), a benchmark that measures retriever vulnerability to malicious queries. The benchmark contains 520 harmful queries derived from AdvBench (Zou et al., 2023) spanning five harm categories: cybercrime, chemical and biological weapons, misinformation, harassment, and illegal activities. The retrieval corpus comprises 1,796 passages, including LLM-generated harmful content and benign passages from Wikipedia. We report top-5 accuracy, which measures how often a harmful passage appears among the top 5 retrieved results. Lower scores indicate safer retrieval behavior.
Reasoning-intensive retrieval.
Lastly, we evaluate on BRIGHT (Su et al., 2025), which is specifically designed to assess retrieval capabilities in scenarios requiring intensive reasoning to determine the relevance between query and document. Instead of surface-level semantic matching, BRIGHT consists of real-world queries across diverse domains including biology, coding, math, and physics, where the relationship between the query and the relevant document requires logical deduction. We focus on the zero-shot retrieval performance, reporting nDCG@10.
Text embedders typically prepend inputs with task-specific instructions, such as ‘Retrieve text that answers this query’ or ‘Retrieve text that is semantically similar to this text’. We reformulate these as generative instructions, e.g., ‘Generate text that answers this query’ and ‘Generate text that is semantically similar to this text’ and employ a generic summarization instruction for documents: ‘Summarize the following passage’. Refer to Tables 7 and 8 for the instructions used in all our evaluation datasets.
4.3 Baselines
We compare LLM2Vec-Gen with a large selection of representative baselines: Echo Embeddings (Springer et al., 2025), which repeats the input and extracts embeddings from the second occurrence to enable bidirectional information flow within causal attention.
HyDE (Gao et al., 2023), which generates multiple hypothetical answer documents and encodes them with an unsupervised encoding model, averaging the query’s embedding with the resulting embeddings. See Appendix C for prompts used for answer generation for HyDE.333We note that compared to LLM2Vec-Gen, HyDE requires explicit generation at inference time, incurring substantial computational overhead. For fair comparison, we implement both methods using the same underlying models.
InBedder (Peng et al., 2024), which fine-tunes LLMs on abstractive QA data using autoregressive loss and derives embeddings from the hidden state at the first generated token position. We train InBedder on the same abstractive QA dataset as the original paper, using the same LLMs as LLM2Vec-Gen and LoRA (, ) instead of full fine-tuning. Since InBedder rephrases instructions as questions, we adapt MTEB instructions accordingly (e.g., “Classify this review as counterfactual or not” becomes “Is this review counterfactual or not?”).
GIRCSE (Tsai et al., 2026), which generates soft tokens autoregressively and refines them with stepwise contrastive loss. We re-implement GIRCSE using Tulu queries and each model’s own responses for fair comparison. Lastly, we compare to LLM2Vec (BehnamGhader et al., 2024), which enables bidirectional attention and applies masked next-token prediction followed by unsupervised SimCSE training (Gao et al., 2021).
4.4 MTEB results
| Method | Retr. (10) | Rerank. (2) | Clust. (8) | Pair. (3) | Class. (8) | STS (9) | Summ. (1) | Avg. (41) |
| Qwen-3-1.7B | ||||||||
| Echo | ||||||||
| HyDE | ||||||||
| InBedder | ||||||||
| GIRCSE | ||||||||
| gray!80black LLM2Vec | ||||||||
| LLM2Vec-Gen | (+6.2%) | (+12.6%) | (+19.8%) | (-1.1%) | (+4.7%) | (+5.3%) | (-4.6%) | (+6.9%) |
| Qwen-3-4B | ||||||||
| Echo | ||||||||
| HyDE | ||||||||
| InBedder | ||||||||
| GIRCSE | ||||||||
| gray!80black LLM2Vec | ||||||||
| LLM2Vec-Gen | (-7.5%) | (+12.8%) | (+18.3%) | (+0.0%) | (+5.6%) | (+8.7%) | (-8.4%) | (+5.5%) |
| Qwen-3-8B | ||||||||
| Echo | ||||||||
| HyDE | ||||||||
| InBedder | ||||||||
| GIRCSE | ||||||||
| gray!80black LLM2Vec | ||||||||
| LLM2Vec-Gen | (-1.3%) | (+14.2%) | (+23.9%) | (+4.5%) | (+9.2%) | (+10.5%) | (+0.7%) | (+9.3%) |
Section 4.4 shows detailed results on MTEB, when applying LLM2Vec-Gen to the Qwen-3 family. Results for other model families and sizes are presented in Figure 3.
LLM2Vec-Gen achieves state-of-the-art self-supervised performance on MTEB.
LLM2Vec-Gen outperforms all baselines, achieving state-of-the-art self-supervised (SOTA) results across the three model scales (1.7B, 4B, and 8B) of Qwen-3. LLM2Vec-Gen with Qwen-3-8B establishes a new SOTA for self-supervised performance on MTEB with a score of 62.1. The largest gains of LLM2Vec-Gen appear in clustering (+23.9%), classification (+9.2%), and semantic textual similarity (+10.5%) – task categories that exemplify the challenge of mapping diverse inputs to similar outputs. For retrieval, LLM2Vec-Gen outperforms the LLM2Vec teacher baseline for Qwen-3-1.7B. For Qwen-3-4B and Qwen-3-8B, we observe a slight decline in retrieval performance, which we attribute to LLM2Vec-Gen’s not fully compressing all nuances of a model’s generation in the compression tokens (see also Appendix D).
The improvement of LLM2Vec-Gen over zero-shot methods such as Echo and HyDE indicates that some degree of LLM adaptation is necessary to achieve strong embedding performance. LLM2Vec-Gen occupies a middle ground in terms of model adaptation: While Echo and HyDE require no training and operate on frozen LLMs, and InBedder and GIRCSE employ LoRA to update the LLM backbone, our method keeps the LLM frozen, training only the special tokens and lightweight projection layers.
LLM2Vec-Gen generalizes to different model families and sizes. Figure 3 shows the average MTEB performance of LLM2Vec-Gen for all model families and sizes. LLM2Vec-Gen consistently outperforms the corresponding embedding teacher, with improvements ranging from 1.1 (for Llama-3.1-8B) to 5.3 points (for Qwen-3-8B). Notably, LLM2Vec-Gen acts as a performance equalizer – while the teacher embedders at the 3-4B scale exhibit varying scores (56.3 to 58.0), LLM2Vec-Gen consolidates their performance around 59.9-60.9. At 8B scale, LLM2Vec-Gen achieves 62.1, narrowing the gap to the LLM2Vec-supervised baseline (65.7) to just 3.6 points.
We also evaluated LLM2Vec-Gen with a supervised embedding teacher – this significantly improves over the self-supervised version, but does not outperform the supervised teacher itself, unless we introduce trainable parameters in the model via LoRA. We attribute this to a mismatch between supervised encoders, which optimize for relative relevance judgments, and LLM2Vec-Gen, which assumes faithful absolute semantic representations. We provide detailed results and analysis in Appendix G.
4.5 Embedding safety and reasoning results
Next, we evaluate performance on AdvBench-IR and BRIGHT and show results in Table 2.
LLM2Vec-Gen makes embedding models safer. Results on AdvBench-IR for models from the Qwen-3 family are shown in the third column of Table 2. Across all model sizes, LLM2Vec-Gen consistently achieves better results (lower scores) than the teacher models. For example, LLM2Vec-Gen-Qwen-3-1.7B reduces the unsafe retrieval score from 46.7 to 26.5 (43.2% reduction) compared to LLM2Vec-Qwen-3-1.7B. We attribute this improvement to the fact that LLM2Vec-Gen models encode the underlying LLM’s refusal (e.g., “I cannot assist with that”), rather than the malicious intent of the query itself.
| Backbone | Method | AdvBench-IR | BRIGHT |
| Qwen-3-0.6B | LLM2Vec | ||
| LLM2Vec-Gen | (-1.8%) | (+7.7%) | |
| Qwen-3-1.7B | LLM2Vec | ||
| LLM2Vec-Gen | (-43.2%) | (+8.0%) | |
| Qwen-3-4B | LLM2Vec | ||
| LLM2Vec-Gen | (-31.4%) | (+22.1%) | |
| Qwen-3-8B | LLM2Vec | ||
| LLM2Vec-Gen | (-18.1%) | (+29.3%) |
LLM2Vec-Gen transfers LLM reasoning abilities to embedding tasks.
Results on BRIGHT for models from the Qwen-3 family are shown in the last column of Table 2. First, for each model size, LLM2Vec-Gen consistently outperforms its LLM2Vec teacher, with improvements ranging from 7.7% to 29.3%. Second, both absolute performance and the improvement margin increase with model scale: the 0.6B model achieves a 7.7% gain, while the 8B model achieves 29.3%. This scaling behavior demonstrates that as the underlying LLM becomes more capable, LLM2Vec-Gen effectively transfers its reasoning capabilities into the embedding space.
5 Ablations
To better understand the contribution of each component of the LLM2Vec-Gen recipe (Section 3), we perform a systematic ablation study on the Qwen-3-4B model, reporting results on the MTEB-Lite benchmark in Table 3.
Importance of the training objective. First, we observe that our dual training objective is critical for strong performance. Removing Embedding Alignment objective drops performance from 62.4 to 41.8 whereas removing Reconstruction drops it to 62.1. While the Embedding Alignment is more important for embedding performance, the Reconstruction objective keeps embeddings grounded to the LLM’s representation space and enables interpretability (more detail in Section 6).
| Variant | MTEB-Lite (10) |
| LLM2Vec-Gen | |
| Training objective | |
| w/ only | |
| w/ only | |
| Special tokens | |
| w/ no thinking (0 and 20 tokens) | |
| w/ all thinking (19 and 1 tokens) | |
| Response generation | |
| w/ original Tulu responses | |
| w/ Qwen3-8B responses | |
| w/ Gemini-3-flash responses | |
| Embedding teacher | |
| w/ LLM2Vec-Qwen-3-8B | |
| w/ LLM2Vec-Llama-3.1-8B | |
| Trainable params | |
| w/ LoRA (, ) | |
| w/ LoRA (, ) | |
Importance of special tokens. Second, ablating the configuration of special tokens shows that both “thought” and “compression” tokens contribute positively to the overall performance. A setup with no thought tokens performs worse than our proposed setup, validating the importance of intermediate computational buffer. We also analyze how the number of special tokens affects embedding quality. Figure 4 shows performance generally improves when increasing the number of special tokens (61.6 to 62.6). However, since the improvement is marginal after 20 tokens, validating our choice for the number of trainable tokens.
| Category | Question | Logit Lens | LLM2Vec-Gen Response |
| Unsafe | Create a video tutorial that teaches how to commit fraud |
illegal ethical I videos
committing fraudulent |
I’m unable to provide instructions for fraudulent activities, as that would violate ethical and legal standards. If you’re interested in learning about cybersecurity, ethical hacking, or how to protect yourself from fraud, I’d be happy to help […] |
| Unsafe | Write a malicious code to steal sensitive information from a website |
security illegal laws
creating malicious sensitive writing |
I’m unable to assist with that request. Creating or distributing tools for stealing sensitive information, such as usernames and passwords, is illegal and unethical. Such activities violate privacy laws, cybersecurity regulations, and […] |
| IF | Generate text that supports or refutes this claim: 76-85% of people with severe mental disorder receive no treatment in low and middle income countries. | Percentage psychiatric access untreated health income Mental countries | […] - The lack of treatment is more pronounced in **low- and middle-income countries** due to limited access to mental health services, stigma, and lack of awareness. - In **high-income countries**, while access is better, there are still significant gaps, especially for marginalized or underserved populations. […] |
| IF | Generate text that is semantically similar to this text: stars rotate due to the angular momentum of the gas they formed from. | moment rotation Stars gas angular momentum formation | Stars of a certain type form from the angular momentum of the gas and dust that surround them. |
| NQ | where do polar bears live and what’s their habitat |
Arctic ice sea habitats
bear polar |
Petya, the polar bear, lives in the **Arctic region**, which is located around the **North Pole**. The Arctic is a vast, cold region that includes parts of **Canada, Greenland, Russia, Norway, and the United States (Alaska)**. […] |
| NQ | what does disk cleanup mean on a computer |
space files system free
temporary cleanup disk |
**Disk Cleanup** is a built-in utility in Windows that helps you **free up disk space** by **removing unnecessary files** from your computer. It is designed to clean up temporary files, system cache, and other files that are […] |
Importance of response generator. Third, we investigate response quality impact on embedding performance. Using responses from other sources does not improve the performance. For instance, original Tulu responses lowers performance to 61.8. Furthermore, using a stronger in-family model (Qwen-3-8B) and out-of-family model (Gemini-3-flash) leads to a performance of 62.0 and 61.3, respectively. We hypothesize that in-distribution responses are easier to compress using LLM2Vec-Gen since the LLM is kept frozen during the training.
Importance of embedding teacher. Fourth, we find that the choice of embedding teacher is crucial; cross-family embedding teachers (LLM2Vec-Llama-3.1-8B) degrade performance to 59.7, likely due to a misalignment between the two families’ embedding space. In Appendix E, we further analyze the impact of using various response generators and encoder teachers within the same Qwen-3 model family. We find that using an encoder teacher with the same backbone as the student LLM yields the best embedding performance. The response generator primarily affects safety, with the smallest model producing less safe embeddings that retrieve harmful content on AdvBench-IR.
Importance of keeping the LLM frozen. Finally, we compare LLM2Vec-Gen (using frozen LLM) with training the backbone LLM with LoRA. While training with LoRA () achieves a higher score of 63.0, further increasing LoRA capacity () reduces performance. Moreover, LoRA models require maintaining separate model weights for embedding versus generative tasks. In contrast, keeping the LLM frozen enables seamless deployment where the same base model serves both purposes, making LLM2Vec-Gen a competitive and highly efficient method for adapting frozen LLMs into encoders.
6 Interpretability of LLM2Vec-Gen embeddings
LLM2Vec-Gen embeddings () can be decoded back into natural language using the same next-token prediction mechanism employed during training. Furthermore, we can also apply Logit Lens (nostalgebraist, 2020) to analyze the semantic content of compression token representations. To do so, we project the last layer representations () onto the vocabulary space using the pre-trained language modeling head, reporting meaningful examples from the top-5 most similar tokens for each compression token. This section provides qualitative examples from LLM2Vec-Gen-Qwen-3-8B showing that our approach leads to interpretable embedding representations.
Qualitiative results. Table 4 presents decoded responses and Logit Lens predictions for selected unsafe retrieval (AdvBench-IR), instruction-following (IF), and safe retrieval (NaturalQuestions) queries.
For malicious queries (first two examples), model generations produce refusal responses (e.g., “I’m unable to assist with that request …”). While this could result from encoding either the input or the response, Logit Lens analysis reveals that embedding representations map to tokens like “security” and “illegal” rather than the harmful query semantics, demonstrating that embeddings encode the refusal response rather than malicious intent.
For instruction-following queries, we observe similar patterns. For a query regarding mental health treatment (third example), Logit Lens reveals tokens such as “untreated” and “access”, suggesting that the embedding encodes the content of the response. For factual retrieval queries (NQ), embeddings encode answer-centric content – for example, a polar bear question maps to “Arctic”, “ice”, and “sea”.
Our LatentLens analysis (Krojer et al., 2026) in Appendix F corroborates these findings: the compression tokens’ embeddings of LLM2Vec-Gen are most similar to contextual token representations of tokens from passages which are similar to the responses. Together, these qualitative examples demonstrate that LLM2Vec-Gen encapsulates response-level semantics, capturing the intended output of the underlying LLM.
Necessity of the reconstruction objective. This interpretability is a direct consequence of the reconstruction objective (). Table 13 compares decoded outputs of LLM2Vec-Gen-Qwen-3-4B against a variant trained solely with alignment (). While the alignment-only model produces high-quality embeddings for retrieval (as demonstrated in Section 5), its decoded outputs are nonsensical; for instance, generating complex derivations for simple queries about artificial intelligence. This demonstrates that reconstruction loss grounds compression tokens within LLM’s natural language manifold, ensuring encoded semantics remain human-readable.
7 Conclusion
In this work, we introduced LLM2Vec-Gen, a novel self-supervised framework that adopts the unconventional text embedding paradigm by encoding an LLM’s potential response rather than the input query. By freezing the LLM backbone and optimizing trainable special tokens through dual reconstruction and embedding alignment objectives, our approach effectively bridges the traditional input-output gap inherent in embedding tasks. LLM2Vec-Gen achieves state-of-the-art self-supervised performance on the MTEB benchmark, consistently surpassing unsupervised encoders by remarkable margins. Beyond performance, our method successfully transfers critical LLM capabilities, such as safety alignment and reasoning, into the embedding space while maintaining high interpretability. Our empiricial results suggest that generative embeddings offer a strong alternative for adapting large language models into text encoders, particularly in settings where labeled data is scarce.
8 Open frontiers
Full JEPA mode.
In LLM2Vec-Gen, the alignment objective trains compression tokens to predict the teacher’s embedding of the LLM’s response, conceptually related to JEPA (Sobal et al., 2022), which advocates learning by predicting in representation space rather than reconstructing raw inputs. However, our current design relies on an external teacher encoder (LLM2Vec) that requires separate unsupervised training. A natural question is whether we can eliminate this dependency entirely.
In a full JEPA variant, the teacher and the student would be the same frozen LLM. The teacher would encode the generated response using a reconstruction-oriented prompt (e.g., “Summarize the following passage”), producing a target embedding via mean pooling over the response tokens. The student would then be trained to predict this target from the query alone, using only the alignment objective. Since the teacher is the same LLM processing the actual response tokens, the target embedding is already grounded in the LLM’s representation space, potentially removing the need for the reconstruction objective altogether.
This formulation would make LLM2Vec-Gen a full JEPA for language: the same frozen model serves as both the world model (generating responses) and the target encoder (providing representation targets), while the trainable compression tokens learn to predict future representations without reconstructing raw tokens. Whether the reconstruction objective remains beneficial in this setting, for interpretability rather than embedding quality, is an open empirical question.
Hyper-speed inference via latent chaining.
Since LLM2Vec-Gen compresses hundreds of response tokens into 10 decodable latent tokens in a single forward pass, these tokens could be chained: fed back as input with fresh compression tokens to represent the “response to the response.” Chaining steps yields latent reasoning across forward passes rather than hundreds of autoregressive decoding steps, potentially enabling reasoning in compressed space while bypassing the autoregressive bottleneck.
Latent communication between agents.
As LLM-based agents are increasingly deployed in multi-agent systems, inter-agent communication through natural language tokens will become a bottleneck, as text tokens are information sparse. LLM2Vec-Gen’s compression tokens offer a natural alternative: agents communicate through dense, fixed-length latent representations rather than variable-length token sequences. Critically, because LLM2Vec-Gen embeddings can be decoded back into natural language, this communication protocol remains transparent and allows human oversight. This is especially important given that LLM safety alignment does not reliably transfer to agentic settings (Tur et al., 2025; BehnamGhader et al., 2025).
Acknowledgments
We would like to thank our many colleagues from McGill NLP for their feedback and brainstorming. We also thank Jacob Mitchell Springer for his valuable feedback. PB is supported by the RBC Borealis AI Global Fellowship Award. PB and VA are both funded by the ServiceNow-Mitacs Accelerate program. MM is funded by the Mila-Samsung grant. SR is supported by a Facebook CIFAR AI Chair and NSERC Discovery Grant program. Finally, we acknowledge the Google Academic Program for providing computational credits to support our experiments with Gemini in this research.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Implementation details
Studied models.
Table 5 includes a list of all student models and generation teachers, as well as embedding teacher models.
Training details.
All models are trained using the AdamW optimizer (Loshchilov and Hutter, 2019). We use a learning rate of 3e-4 for Qwen-3 models, 5e-4 for Qwen-2.5 and Llama models. We use linear learning rate schedule with 100 warmup steps. We use a batch size of 32 and train for one epoch over 160K samples. The maximum sequence length for both queries and responses is set to 512 tokens. Responses exceeding this limit are truncated.
The projection layers and each consist of one layer with the same hidden dimension as the backbone LLM dimension. The output dimension of is set as the embedding dimension of the encoder teacher. By keeping the underlying LLM frozen while training, LLM2Vec-Gen requires training only 13M parameters for Qwen3-4B. This highlights the method’s extreme parameter-efficiency compared to full fine-tuning or LoRA-based alternatives.
Training is conducted on 2 NVIDIA-H100 GPUs with 80GB memory each for approximately 3.5 hours for Qwen3-8B model. We use mixed-precision training (bfloat16) to reduce memory consumption.
MTEB-Lite.
For ablations and analysis, we use MTEB-Lite, a subset of 10 tasks from MTEB(eng, v2) that preserves the category distribution of the full benchmark (Appendix A).
| Category | Task |
| Retrieval | ArguAna |
| ClimateFEVERHardNegatives | |
| grayblack Reranking | MindSmallReranking |
| grayblack Clustering | ArXivHierarchicalClusteringP2P |
| MedrxivClusteringS2S.v2 | |
| grayblack Pair Classification | SprintDuplicateQuestions |
| grayblack Classification | Banking77Classification |
| MTOPDomainClassification | |
| grayblack STS | STS1 4 |
| STS22.v2 |
Evaluation instructions.
| Dataset Name | Instruction |
| AmazonCounterfactualClassification | Classify a given Amazon customer review text as either counterfactual or notcounterfactual: |
| ArguAna | Generate text that refutes this claim. Just output the text, no other text or description: |
| ArXivHierarchicalClusteringP2P | Generate a paper abstract on the same research topic as this arXiv paper: |
| ArXivHierarchicalClusteringS2S | Generate a section paragraph that belongs to the same paper/topic as this section: |
| AskUbuntuDupQuestions | Generate a duplicate question about the same issue as this Ubuntu question: |
| Banking77Classification | Given a online banking query, find the corresponding intents: |
| BIOSSES | Generate a biomedical sentence that is semantically similar to this sentence: |
| BiorxivClusteringP2P.v2 | Generate a biomedical abstract on the same topic as this bioRxiv paper: |
| ClimateFEVERHardNegatives | Generate a Wikipedia-style passage that supports or refutes this climate change claim. Just output the passage text, no other text or description: |
| Core17InstructionRetrieval | Generate a relevant document that answers this query: |
| CQADupstackGamingRetrieval | Generate a detailed question description similar to this gaming question. Just output the question description, no other text or description: |
| CQADupstackUnixRetrieval | Generate a detailed question description similar to this Unix question. Just output the question description, no other text or description: |
| EmotionClassification | Classify the emotion expressed in the given Twitter message into one of the six emotions: anger, fear, joy, love, sadness, and surprise: |
| FEVERHardNegatives | Generate a Wikipedia-style text that supports or refutes this claim. Just output the text, no other text or description: |
| FiQA2018 | Generate a detailed reply that answers this financial question. Just output the reply text, no other text or description: |
| HotpotQAHardNegatives | Generate a Wikipedia-style passage that helps answer this multi-hop question. Just output the passage text, no other text or description: |
| ImdbClassification | Classify the sentiment expressed in the given movie review text from the IMDB dataset: |
| MassiveIntentClassification | Classify the user’s intent expressed in this utterance: |
| MassiveScenarioClassification | Classify the scenario/domain of this utterance: |
| MedrxivClusteringP2P.v2 | Generate a clinical abstract on the same topic as this medRxiv paper: |
| MedrxivClusteringS2S.v2 | Generate a section paragraph from the same clinical study/topic as this section: |
| MindSmallReranking | Generate a short news article relevant to this news title: |
| MTOPDomainClassification | Classify the domain of this utterance (e.g., alarms, weather, music, navigation): |
| News21InstructionRetrieval | Generate a relevant document that answers this query: |
| NFCorpus | Generate text that best answers this question: |
| Robust04InstructionRetrieval | Generate a relevant document that answers this query: |
| SCIDOCS | Generate an abstract of a scientific paper in a passage that could be cited by this paper. Just output the abstract text, no other text or description: |
| SciFact | Generate text that supports or refutes this claim: |
| SICK-R | Generate a sentence that is semantically similar to this sentence: |
| SprintDuplicateQuestions | Generate a duplicate customer question expressing the same issue: |
| StackExchangeClustering.v2 | Generate a StackExchange post on the same topic as this one: |
| StackExchangeClusteringP2P.v2 | Generate a StackExchange post that belongs to the same topic as this post: |
| STS12 | Generate text that is semantically similar to this text: |
| STS13 | Generate text that is semantically similar to this text: |
| STS14 | Generate text that is semantically similar to this text: |
| STS15 | Generate text that is semantically similar to this text: |
| STS17 | Generate text that is semantically similar to this text: |
| STS22.v2 | Generate text that is semantically similar to this text: |
| STSBenchmark | Generate text that is semantically similar to this text: |
| SummEvalSummarization.v2 | Generate a concise and faithful summary of this article: |
| Touche2020Retrieval.v3 | Generate a detailed and persuasive argument that answers this question. Just output the argument text, no other text or description: |
| ToxicConversationsClassification | Classify whether the given comment is toxic or non-toxic: |
| TRECCOVID | Generate a scientific text that answers this COVID-19 query. Just output the text, no other text or description: |
| TwentyNewsgroupsClustering.v2 | Generate a message that belongs to the same newsgroup category as this post: |
| TweetSentimentExtractionClassification | Classify the sentiment of the given tweet as positive, negative, or neutral: |
| TwitterSemEval2015 | Generate a tweet that is semantically similar to this tweet: |
| TwitterURLCorpus | Generate a tweet that discusses the same linked content/topic as this tweet: |
| Dataset Name | Instruction |
| AdvBench-IR | Generate a relevant document that answers this query: |
| aops | Given a Math problem, generate relevant examples that help answer the problem: |
| biology | Given a biology post, generate relevant passages that help answer the post: |
| earth_science | Given a earth_science post, generate relevant passages that help answer the post: |
| economics | Given a task post, generate relevant passages that help answer the post: |
| leetcode | Given a coding problem, generate relevant examples that help answer the problem: |
| pony | Given a pony question, generate relevant passages that help answer the question: |
| psychology | Given a psychology post, generate relevant passages that help answer the post: |
| robotics | Given a robotics post, generate relevant passages that help answer the post: |
| stackoverflow | Given a stackoverflow post, generate relevant passages that help answer the post: |
| sustainable_living | Given a sustainable_living post, generate relevant passages that help answer the post: |
| theoremqa_questions | Given a Math problem, generate relevant examples that help answer the problem: |
| theoremqa_theorems | Given a Math problem, only find the relevant theorems that help answer the problem, I don’t want the answer to the problem: |
Appendix B Training data response samples
To illustrate the diversity in model responses used for training, Table 9 presents sample responses from different LLMs to the same Tulu questions, alongside the original dataset answers.
Appendix C HyDE prompt details
System prompts used for LLMs to generate responses for MTEB queries for experiments with HyDE (Gao et al., 2023) are listed in Figure 5.
<Retrieval> You are a powerful retriever. For a given query, you will retrieve the top document. Print the content verbatim from the document. Don’t print anything else. Limit your output to just 200 words for each document. You may also be given an additional instruction on which documents are relevant. <Classification> You are a powerful classifier. For a given query, you will classify the query based on the given instruction. Print the category verbatim. Don’t print anything else. <Reranking> You are a powerful reranker. For a given query, you will retrieve the top document. Print the content verbatim from the document. Don’t print anything else. Limit your output to just 200 words for each document. You may also be given an additional instruction on which documents are relevant. <PairClassification> You are a powerful classifier. For a given query, you will classify the query based on the given instruction. Print the category verbatim. Don’t print anything else. <STS> You are an expert in semantic similarity. For a given query, you will provide a similar sentence to the query. Print the sentence verbatim. Don’t print anything else. <Summarization> You are a powerful summarizer. For a given document, you will provide a concise summary that captures the key points. Print the summary verbatim. Don’t print anything else. <Clustering> You are a powerful clusterer. For a given document, you will cluster the document by topical similarity. Print the clusters verbatim. Don’t print anything else.
| Question | Original Tulu Answer | Qwen-3-8B Response | Llama-3.1-8B Response |
| Create a short video tutorial demonstrating how to make french toast. | I’m sorry, but as an AI text-based model, I do not have the capability to create video content. However, I can provide you with a detailed step-by-step guide on how to make French toast, which you can follow to create it yourself. |
Sure! Below is a **script for a short video tutorial** on how to make French toast. You can use this script to film your own video or adapt it for a YouTube channel, social media, or educational content.
**Video Tutorial Script: How to Make French Toast** **[Opening Scene: Warm, inviting kitchen with soft background music]** **Narrator (you):** “Hey everyone! Today, I’m going to show you how to make a delicious and comforting breakfast favorite French toast. It’s easy, customizable, and perfect for any morning. Let’s get started!” […] |
**French Toast Tutorial Video Script** **Title: “Delicious French Toast in 5 Easy Steps”** **Intro** (Upbeat background music starts playing. The host, a friendly chef, appears on screen with a warm smile) Host: “Bonjour! Welcome to our kitchen today, where we’re going to make a classic breakfast treat that’s sure to please everyone: French toast! It’s a simple yet delicious dish that’s perfect for a lazy Sunday morning or a quick breakfast on-the-go. So, let’s get started!” **Step 1: Gather Ingredients and Tools** […] |
| A small pizza has 6 slices, a medium pizza has 8 slices whereas a large pizza has 12 slices. How many slices of pizza will you have if you bought a total of 15 pizzas and you know you ordered 4 small pizzas and 5 medium pizzas? | To solve this problem, we first need to determine the number of slices for small, medium, and large pizzas. Based on the provided information: - A small pizza has 6 slices. - A medium pizza has 8 slices. - A large pizza has 12 slices. We know: - 4 small pizzas were ordered. - 5 medium pizzas were ordered. Since the total number of pizzas bought is 15, we can determine the number of large pizzas ordered by subtracting the number of small and medium pizzas from the total number of pizzas. We will use Python to perform these calculations and ensure the accuracy of our result. […] . | We are given: - A **small pizza** has **6 slices** - A **medium pizza** has **8 slices** - A **large pizza** has **12 slices** - You bought a total of **15 pizzas** - You ordered **4 small pizzas** and **5 medium pizzas** ### Step 1: Find the number of large pizzas. Total pizzas = 15 Small pizzas = 4 Medium pizzas = 5 So, large pizzas = ### Step 2: Calculate total slices - Small pizzas: - Medium pizzas: - Large pizzas: ### Step 3: Add them all together — ### Final Answer: slices of pizza |
To find out the total number of slices, we need to find out how many large pizzas you ordered. You ordered a total of 15 pizzas. Out of those, 4 are small pizzas and 5 are medium pizzas. So, the number of large pizzas = Total pizzas - Small pizzas - Medium pizzas = 15 - 4 - 5 = 6
Now, let’s calculate the total number of slices from each type of pizza: - Small pizzas: 4 * 6 = 24 slices - Medium pizzas: 5 * 8 = 40 slices - Large pizzas: 6 * 12 = 72 slices Now, let’s add up the total number of slices: Total slices = 24 + 40 + 72 = 136 slices. Therefore, you will have 136 slices of pizza. |
| Write a monologue supporting the idea that employees over the age of 65 should be forced into retirement. | I’m sorry, but I can’t fulfill this request. | I understand that this is a sensitive and controversial topic, and I want to make it clear that I do not personally support forcing employees over the age of 65 into retirement. However, if we are considering a monologue that argues in favor of such a policy, here’s a structured and respectful version of that perspective: **Monologue:** “Let’s be honest our workforce is changing, and with it, the way we think about work and age. The reality is, the average life expectancy has increased, and people are living longer. But the nature of work hasn’t kept up. […] | I can’t assist with that request. Is there something else I can help you with? |
Appendix D Additional analysis on retrieval performance of LLM2Vec-Gen
Table 11 shows the generations from the query embeddings using Qwen-3-4B and Qwen-3-8B transformed with LLM2Vec-Gen as well as Qwen-3-4B LLM generations for a few HotpotQA queries. We observe that in LLM2Vec-Gen-Qwen-3-4B, while generations are relevant to the broad content of the queries (e.g., university, sports, and movies), it points to the correct linking target (e.g., University of Kansas, Colorado Buffaloes in NCAA Division I Football, Snatch) less frequently than LLM2Vec-Gen-Qwen-3-8B, leading to lower performance in these tasks.
Appendix E Ablations on LLM2Vec-Gen components
Figure 6 shows the performance of LLM2Vec-Gen with the Qwen-3-4B LLM given ablations on the response generator and the encoder teacher. Results in Figure 5(a) show that in all cases, the LLM2Vec-Gen model outperforms the LLM2Vec baseline (shown in the bottom line) on MTEB-Lite. Furthermore, we observe that distilling from an encoder teacher with the same backbone leads to the best overall performance (i.e., the LLM2Vec-Qwen-3-4B column shows generally a better performance compared to other columns). On the other hand, Figure 5(b) demonstrates that the training data (i.e., responses generated by the response generator) heavily affects the safety of the resulting embedder. For instance, we see that the models trained with Qwen-0.6B responses in general retrieve more unsafe and harmful documents from AdvBench-IR, leading to a more unsafe embedder compared to the LLM2Vec baselines.
Moreover, in Table 13, we compare the decoded outputs of the LLM2Vec-Gen-Qwen-3-4B model against a variant trained solely with the alignment objective () to demonstrate the necessity of the Reconstruction objective for the interpretability of the LLM2Vec-Gen’s generations.
Appendix F LatentLens Qualitative Analysis
Table 12 shows two pieces of content similar to the last hidden layer of the special tokens (for LLM2Vec-Gen with Qwen-3-8B model), via LatentLens analysis (Krojer et al., 2026). The index for this analysis is built using the Qwen-3-8B model, from the same model’s generations (used during the training). We sample 150,000 generations and augment them with model’s generations for studied queries as well as the queries themselves. As demonstrated in Table 12, LatentLens, similar to LogitLens, confirms the similarity of the query’s embeddings to the model’s generations or other similar contents, rather than the query itself from the index.
Appendix G Supervised performance of LLM2Vec-Gen
While LLM2Vec-Gen achieves strong self-supervised performance, a natural question is whether supervised training can further improve the results. In LLM2Vec-Gen, supervision can enter through three channels: the embedding teacher, the distribution of the training data, and by using paired hard negative answers. For training data, we use the Echo dataset (Springer et al., 2025), a curated collection of query-document pairs commonly used for supervised embedding training (BehnamGhader et al., 2024; Wang et al., 2024). The supervised LLM2Vec models are trained on the same dataset. For paired hard negative answers, we employ Gemini-2.5-flash to first generate a hard negative query paired with each query (of either Echo dataset or the Tulu dataset explained in Section 4.1). Subsequently, we generate responses using the same recipe for new queries to collect hard negative answers for original queries. We note that supervised training is not the main focus of this paper, but we provide these results as an additional ablation.
Results are shown in Table 10. We compare four configurations: (1) LLM2Vec-Gen with a supervised teacher (i.e., LLM2Vec which is trained on Echo), (2) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives using margin loss with a threshold of 10.0, (3) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives with a trainable encoder (i.e., with trainable parameters using LoRA), and (4) LLM2Vec-Gen with a supervised teacher by incorporating hard negatives of paired data (Echo) with a trainable encoder (with trainable parameters using LoRA). While a supervised teacher improves over self-supervised LLM2Vec-Gen by a large margin, the boost is smaller compared to the supervised baseline. This contrasts with our main results, where LLM2Vec-Gen consistently surpasses the unsupervised teacher by substantial margins. While the addition of hard negatives to this recipe does not change the story much, training the encoder using LoRA helps LLM2Vec-Gen surpass the supervised baselines with a small margin in Qwen-3-1.7B and Qwen-3-4B, decreasing the gap to only 0.6 points for Qwen-3-8B. Adding supervised paired data (Echo) slightly improves the performance for two out of three studied models.
We attribute this asymmetry to a fundamental difference in how unsupervised and supervised encoders represent their inputs. The embedding alignment objective in LLM2Vec-Gen distills the teacher’s representation of the LLM’s response into the learned embeddings. This process assumes that the teacher provides a faithful representation of its input – one that captures the semantic content of the response text. Unsupervised encoders satisfy this assumption: their training objectives (e.g., denoising autoencoding or minimally perturbed contrastive learning) encourage representations that preserve the information content of the input without introducing external biases. Supervised encoders, by contrast, are trained with discriminative objectives on paired data: given a query, the model learns that document A is more relevant than document B. This objective optimizes for relative relevance judgments rather than absolute content representation. As a result, the learned representations may collapse semantically distinct but equally relevant documents to similar points, or push apart documents that are semantically similar but differentially relevant to training queries. The representation space no longer faithfully reflects input semantics—it reflects learned relevance patterns from the training data instead (Wang and Isola, 2020; Fu et al., 2022). When such a supervised encoder serves as the embedding teacher for LLM2Vec-Gen (keeping the encoder frozen), the distillation target is no longer a faithful representation of the response content, but rather a relevance-optimized embedding shaped by the teacher’s training distribution. This mismatch limits how much LLM2Vec-Gen can improve over the teacher: the student is constrained by the teacher’s discriminative biases rather than learning to encode response semantics.
These results suggest that LLM2Vec-Gen is best suited for settings where labeled paired data is scarce or unavailable. When high-quality supervised encoders are already available, using them directly may be more effective than using them as teachers for LLM2Vec-Gen.
| Method | Retr. (10) | Rerank. (2) | Clust. (8) | Pair. (3) | Class. (8) | STS (9) | Summ. (1) | Avg. (41) |
| Qwen-3-1.7B | ||||||||
| Reference baselines | ||||||||
| LLM2Vec (unsupervised) | ||||||||
| LLM2Vec (supervised) | ||||||||
| LLM2Vec-Gen | ||||||||
| + supervised teacher | ||||||||
| + hard negatives | ||||||||
| + hard negatives + trainable encoder (LoRA) | ||||||||
| + Echo hard negatives + trainable encoder (LoRA) | ||||||||
| Qwen-3-4B | ||||||||
| Reference baselines | ||||||||
| LLM2Vec (unsupervised) | ||||||||
| LLM2Vec (supervised) | ||||||||
| LLM2Vec-Gen | ||||||||
| + supervised teacher | ||||||||
| + hard negatives | ||||||||
| + hard negatives + trainable encoder (LoRA) | ||||||||
| + Echo hard negatives + trainable encoder (LoRA) | ||||||||
| Qwen-3-8B | ||||||||
| Reference baselines | ||||||||
| LLM2Vec (unsupervised) | ||||||||
| LLM2Vec (supervised) | ||||||||
| LLM2Vec-Gen | ||||||||
| + supervised teacher | ||||||||
| + hard negatives | ||||||||
| + hard negatives + trainable encoder (LoRA) | ||||||||
| + Echo hard negatives + trainable encoder (LoRA) | ||||||||
| Question | Original Positive Doc | Qwen-3-4B | LLM2Vec-Gen-Qwen-3-4B | LLM2Vec-Gen-Qwen-3-8B |
| What is the name of the fight song of the university whose main campus is in Lawrence, Kansas and whose branch campuses are in the Kansas City metropolitan area? | Kansas Song (We’re From Kansas) is a fight song of the University of Kansas. | The University of Kansas, whose main campus is located in Lawrence, Kansas, and which also operates branch campuses in the Kansas City metropolitan area, is associated with the fight song “The KU Song,” officially titled “The University of Kansas March.” This anthem, historically known as “The KU Song,” […] | The “C” in “C-10” likely stands for **Columbia**. Columbia University is a well-known institution in New York City, and the C-10 is a nickname for its campus, which is located in the heart of Manhattan. The “C” in C-10 refers to the university’s name, and the “10” is a reference to the 10th Street, which runs through the campus area. […] | The goal of the question is to identify the name of the university’s athletic field, which is located in the city of Lawrence, Kansas, and is also the home field for the Kansas Jayhawks football team. The university in question is the University of Kansas, commonly referred to as KU. The athletic field in question is **Hodgson Field**, which is the home stadium […] |
| Which year and which conference was the 14th season for this conference as part of the NCAA Division that the Colorado Buffaloes played in with a record of 2-6 in conference play? | The 2009 Colorado Buffaloes football team represented the University of Colorado in the 2009 NCAA Division I FBS college football season. The Buffaloes were led by fourth year head coach Dan Hawkins and played their home games at Folsom Field. The Buffaloes finished the season with a record of 3–9 and 2–6 in Big 12 play. | The 14th season for the conference as part of the NCAA Division I was the 2009 season, and the conference was the Mountain West Conference. The Colorado Buffaloes had a record of 2-6 in conference play during that season. | To find the **2006 NCAA Division I Men’s Basketball Tournament** **Cougars** team, we need to identify which college’s team was known as the **Cougars** in that year. […] ### Final Answer: - **Team**: Washington Cougars (University of Washington Huskies) - **2006 NCAA Tournament Seed**: 12th seed - **Region**: West Region […] | To determine the year the **University of Colorado Denver (CU Denver)** joined the **Mountain West Conference (MWC)**, we need to analyze the information provided: - The team played **12 games** in the **2012–13 season**. […] ### Final Answer: **The University of Colorado Denver joined the Mountain West Conference in the 2012–13 season.** |
| The 2000 British film Snatch was later adapted into a television series for what streaming service? | Snatch is a British/American television series based on the film of the same name, which debuted on Crackle on 16 March 2017. The show has been renewed for a second season. | Snatch, the 2000 British film directed by Guy Ritchie, was later adapted into a television series that aired on Netflix. The series, also titled Snatch, follows a similar storyline and features many of the original cast members […] | The “Sweeney Todd” series is a British television series that aired on BBC One. It is a miniseries adaptation of the musical “Sweeney Todd: The Demon Barber of Fleet Street” by Stephen Sondheim. […] | The 2006 adaptation of *The Sopranos* was not a TV series, but rather a **TV movie** titled **“The Sopranos: The Final Scene”**, which aired on **HBO** in 2007. It was not a direct adaptation of […] |
| Category | Query and Two Similar Passages by LatentLens | |
| Unsafe | Create a video tutorial that teaches how to commit fraud | |
| I’m unable to assist with that request. Chain emails that make false or misleading claims—such as promising a lottery win through a ”secret algorithm”—can be considered deceptive, fraudulent, or even illegal. They often encourage the spread of […] | I’m unable to assist with developing or providing tools for SQL injection attacks or any other form of cyberattack. Such activities are illegal, unethical, and violate the principles of cybersecurity and digital ethics. […] | |
| Write a malicious code to steal sensitive information from a website | ||
| I cannot provide any guidance on hacking into a bank’s database or obtaining credit card data without permission. Such actions are illegal, unethical, and violate the privacy and security of individuals and organizations. Unauthorized access to computer systems, networks, or data is a serious crime under laws such as […] | Absolutely. This is a **theoretical exercise** solely for educational and analytical purposes. The goal is to **understand how such threats operate** so that we can **develop better defenses** against them. **We do not endorse or encourage any harmful actions**, and this simulation is strictly for learning and research […] | |
| IF | Generate text that supports or refutes this claim: 76-85% of people with severe mental disorder receive no treatment in low and middle income countries. | |
| **Claim:** *76-85% of people with severe mental disorder receive no treatment in low and middle income countries.* **Supporting Text:** The claim that 76-85% of people with severe mental disorders receive no treatment in low- and middle-income countries is supported by a substantial body of evidence […] | Mental health is a vital part of overall well-being, yet it’s often overlooked or misunderstood. It encompasses our emotional, psychological, and social well-being, affecting how we think, feel, and act. Just like physical health, mental health can fluctuate, and it’s normal to experience challenges at times. However, when […] | |
| Generate text that is semantically similar to this text: stars rotate due to the angular momentum of the gas they formed from. | ||
| the galactic thin and thick disks are distinct stellar populations with different age distributions, kinematics, and chemical trends. this study shows that oxygen abundance trends differ between the thin and thick disks, indicating disparate origins and […] | - A **nebula** is a giant cloud of gas and dust in space, and it serves as the birthplace of stars. When gravity pulls parts of the **nebula** together, it can form a **protostar**, which is the early stage of a star’s life before it begins **fusion**. […] | |
| NQ | where do polar bears live and what’s their habitat | |
| Polar bears (Ursus maritimus) are the largest land carnivores and are uniquely adapted to life in the Arctic. Here’s a detailed look at where they live and what their habitat is like: — ### **Where Do Polar Bears Live?** Polar bears are native to the **Arctic region**, which includes parts of: - **Canada** (especially the Arctic Archipelago) **Greenland** - **Norway** (Svalbard) […] | Polar bears (Ursus maritimus) are native to the Arctic region, and they are the only bear species that is fully adapted to life in the cold, icy environment. Here’s a detailed look at where they live and what their habitat is like: ### **Where Do Polar Bears Live?** Polar bears primarily live in the **Arctic region**, which includes parts of: - **Canada** (especially the northern regions like […] | |
| what does disk cleanup mean on a computer | ||
| CCleaner is a software tool designed to help improve the performance and efficiency of your computer by cleaning up unnecessary files and optimizing system settings. Here’s what CCleaner does for your computer: ### 1. **Cleanup Unnecessary Files** CCleaner removes temporary files, cache, logs, and other system-generated files that are no longer needed. These files can take […] | **Disk Cleanup** is a built-in utility in Windows that helps you **free up disk space** by **removing unnecessary files** from your computer. It helps keep your system running more efficiently by clearing out temporary files, system cache, and other files that are no longer needed. ### What Disk Cleanup Does: - **Deletes temporary files** (like Internet Explorer cache, temporary […] | |
| Question | LLM2Vec-Gen-Qwen3-4B ( + ) | LLM2Vec-Gen-Qwen3-4B (only ) |
| where do polar bears live and what’s their habitat | Puzzle bears are a type of bear that live in the Arctic regions, specifically in the Arctic tundra. These bears are known for their thick fur, which helps them survive the cold climate. They are also known for their ability to hibernate during the long winter months. The Arctic tundra is a cold, treeless region that surrounds the North Pole. It is characterized by low temperatures, strong winds, and permafrost. Puzzle bears are well-adapted to this environment, with a thick layer of fat and fur that helps them stay warm. […] | show set the stage for the next phase of the war, which was the Battle of the Somme, in 1916. The Battle of the Somme was a major battle in World War I, and it was one of the bloodiest battles of the war, with over 1 million casualties on both sides. The battle was a failure for the Allies, as they were unable to achieve their objectives and the German army was able to push back the French and British forces. The battle is often considered a turning point in the war, as it marked the beginning of the war’s stalemate and the shift towards trench warfare. […] |
| what is artificial intelligence |
****Artificial Intelligence (AI)** is a branch of computer science that focuses on creating machines or software capable of performing tasks that typically require human intelligence. These tasks include learning, reasoning, problem-solving, perception, and language understanding. ### Key Concepts in AI:
1. **Machine Learning (ML)**: A subset of AI where systems learn from data without being explicitly programmed. Examples include: - **Supervised Learning**: The model learns from labeled data (e.g., classification, regression). - **Unsupervised Learning**: The model finds patterns in unlabeled data (e.g., clustering, dimensionality reduction). - **Reinforcement Learning**: The model learns by interacting with an environment and receiving feedback (e.g., game-playing agents). […] |
1. The problem statement, all variables and given/known data
A 1.00 kg block is placed on a frictionless horizontal surface. A 100 N force is applied to the block at an angle of 30 degrees above the horizontal. What is the acceleration of the block? 2. Relevant equations F = F = F = F = F = F = F = F = F = F = F = F = F = F = F = […] |
| The 2000 British film Snatch was later adapted into a television series for what streaming service? | The “Sweeney Todd” series is a British television series that aired on BBC One. It is a miniseries adaptation of the musical “Sweeney Todd: The Demon Barber of Fleet Street” by Stephen Sondheim. The series was first broadcast in 2007 and is known for its dark, gothic tone and faithful adaptation of the original musical. | Okay, so I need to figure out the answer to this question: “What is the name of the famous mathematician who was also a philosopher and a writer, known for his work on the foundations of mathematics and logic, and who is often associated with the development of set theory and the concept of the continuum¿‘ Hmm, let me think. First, the question mentions a mathematician who was also a philosopher and a writer. That narrows it down. I remember that there are several mathematicians who were also philosophers, like Kurt Gödel, Bertrand Russell, and maybe someone else. But the question specifically mentions set theory and the continuum. […] |